OpenStudy (anonymous):

For what values of n is n! less than 2^n? Justify your answer.

jhonyy9 (jhonyy9):

n! = 1*2*3*.......*n yes ? 2^n= 2^1 or 2^2 or 2^3 or ..... or 2^n yes ? so then you try calculi 1*2=2 2^2 =4 so 2<4 1*2*3 =6 2^3 = 8 so 6<8 1*2*3*4=24 2^4 = 16 so now 24>16

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

yup got it thanks :D